OpenStudy (anonymous):

what is the solution to the inequality below? x^2<36

OpenStudy (anonymous):

choices?

OpenStudy (anonymous):

\[the question is x^{2} \le 36\]

OpenStudy (anonymous):

A. -6\[\le x \le 6\]

OpenStudy (anonymous):

-6 < x < 6

OpenStudy (anonymous):

B. x <-6 or x > 6 C. x\[\le -6 or x \ge 6\] D. -6 < x <6

OpenStudy (anonymous):

Its D